OpenStudy (anonymous):

How do I solve for b^1? a= (b^1+ b^2)/(2) *h. Explain please.

OpenStudy (anonymous):

2*(a/h)-b^2 =b^1

OpenStudy (anonymous):

look at what i did and compare see which is the correct answer 2a/h - b2 = b1 what is the answer?